OpenStudy (anonymous):

Would this be the dimensional analysis of this equation? F=m(V^2/R) (kg)(m/s)^2=(kg)((m/s)^2/(m)) (kg)(m/s)^2=(kg)(m/s)^2 [F]=[MLT-2]? did I do that dimensional analysis right....?

OpenStudy (anonymous):

\[\frac{kg*m}{s^2}=\frac{kg\frac{m^2}{s^2}}{m}=\frac{kg*m^2}{s^2*m}=\frac{kg*m}{s^2}\]

OpenStudy (anonymous):

? Can you draw that out.. It's all mumble jumble on my laptop.. I don't know why.

OpenStudy (anonymous):